OpenStudy (anonymous):

The equation of line WX is y = -2x - 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (-1, -2).

OpenStudy (anonymous):

Select one: a. y = -2x b.y=-1/2x +3/2 c.1/2x - 3/2 d.y=2x

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

i think y= 1/2x -3/2

OpenStudy (nurali):

click best response.